-A chi-square test for goodness-of-fit has a sample size of 50. What are the degrees of freedom for this chi square?

A. 25

B. The degrees of freedom cannot be determined from the information provided.

C. 50

D. 49

-Rodney wants to test the relationship between college graduation rank and annual income. If income is measured on a ratio scale, the appropriate relationship test for Rodney to use is the:

A. chi square test of independence

B. independent-samples t test

C. Spearman r

D. Pearson r

-Sharon is interested in testing the differences between student populations at two different universities. If she randomly selects students from both universities, she is examining _____ sample(s).

single

dependent

independent

confounded
Answer #1

Answer---->>

Que.1)

correct option is D). 49

we know that degrees of freedom for chi-square test for goodness of fit is  n -1 , n = sample size .

Here , given that sample size = n = 50

Therefore , degrees of freedom = n - 1 = 50 - 1 = 49

Que.2)

Given that , Rodney wants to test the relationship between College graduation ranks and income.of income measured on ratio scale. The appropriate relationship test Rodney to use the

  A) chi-square test of independence.

Que.3)

Sharon is interested in testing the differences between Student population at two different universities. If she randomly select students frim both Universities, she examining samples

C) independent.

​​